OpenStudy (anonymous):

Let g be a function that is defined for all x, x≠2, such that g(3) n=4 and the derivative of g is g′(x)= x2–16 x−2 , with x ≠ 2. Find all values of x where the graph of g has a critical value. For each critical value, state whether the graph of g has a local maximum, local minimum or neither. U must justify your answers with a complete sentence. On what intervals is the graph of g concave down? Justify your answer. Write an equation for the tangent line to the graph of g at the point where x = 3. Does this tangent line lie above or below the graph at this point? Justify your answer

OpenStudy (anonymous):

@eliassaab Please helpp!!! will give out medal :)

OpenStudy (anonymous):

@undeadknight26 Please helpp!!! will give out medal :)

OpenStudy (triciaal):

such that g(3) n=4 what does this mean?

OpenStudy (triciaal):

is this what you mean (x2–16) /(x−2)? so that x =4, -4

OpenStudy (triciaal):

g' would be the slope when the slope = 0 then x^2 - 16 = 0 to give those values of x already given that x not = 2 (we will not divide by zero)

OpenStudy (triciaal):

one way to check for maximum or minimum is to get the 2nd derivative at that critical point positive means minimum and negative for maximum

OpenStudy (triciaal):

you can apply the quotient rule to get the derivative

OpenStudy (triciaal):

Write an equation for the tangent line to the graph of g at the point where x = 3. the slope of the tangent line = g' when x = 3 ; replace x with 3 in the equation and compute y - g(3) = -7(x - 3) g(3) n=4 if the n is an error and g(3) = 4 then y - 4= -7(x-3) y= -7x + 21 + 4 tangent line at x = 3 y = -7x + 25

OpenStudy (anonymous):

Yes i'm so sorry I meant g(3)=4 the n is an error @triciaal

OpenStudy (anonymous):

so can you please explain it from the beginning i'm a little confused because of the n that i put (:

OpenStudy (anonymous):

so how do i find all values of x where the graph of g has a critical value? :D @triciaal

OpenStudy (triciaal):

please review all my entries. I assumed it was an error later. I went through the whole problem! what specific question do you have?

OpenStudy (campbell_st):

@triciaal, you method makes sense to me...

OpenStudy (triciaal):

great. thanks.

OpenStudy (anonymous):

awesome! i understood the explanation :) thanks so much i just had to read it over a couple of times doers this tangent line lie above or below the graph at this point? :) Justify the answer

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

i know i have local minimums for both 4 and -4, and an asymptote at x=2. my intervals for the graph are: to the left--> -infinity<x<2 and to the right---> 2>x>infinity

OpenStudy (anonymous):

I think all i have to do is plug in values for x and see if it goes above or below the graph, correct ? :)

OpenStudy (anonymous):

if taht's the case then i think it lies above

OpenStudy (triciaal):

you have a zero at x = 4

OpenStudy (anonymous):

no u have a -3 -7(4)+25 -28+25 = -3

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (triciaal):

@Directrix

OpenStudy (anonymous):

ohh its below

OpenStudy (anonymous):

i think im not sure :P

OpenStudy (triciaal):

did you graph the functions?